2 ∫ 1 (x 2 – 1) dx

21 (x2 – 1) dx

Grade:12

1 Answers

Anusha
9 years ago
Ans:- 21(x2 -1)dx=[(x2+1 /2+1)-x]12 =[(x3/3)-x]12
=[(8/3)-(1/3)]-(2-1)
=(7/3)-1
=(7-3)/3
=4/3